OpenStudy (anonymous):

determine which ordered pairs are solutions for y = 5x -4 A. (0,-4) B. 5, -21) C. (0,-5) D. (4/5,0)

OpenStudy (precal):

recall the points given are (x,y) so (0,-4) is x=0 and y=-4 y=5x-4 -4=5(0)-4 -4=-4 True so therefore A is your solution.

OpenStudy (anonymous):

A. (0,-4), D. (4/5,0)

OpenStudy (anonymous):

x and y intercept points